# A Survey on Cross-Domain Sequential Recommendation

### 10 Jan 2024 v1

@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
Utilizing an overall structural framework for CDSR modeling, we take a more tangible perspective to summarize the technologies adopted by existing models in addressing the challenges. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively.
@@ 4.1 Basic Technologies
According to Section 2.3, the cross-domain information is always considered as a pure sequence representation and a graph representation. So we illustrate the utilization of sequence modeling technologies (i.e., recurrent neural networks and attention mechanisms) and graph structure modeling technologies (i.e., graph neural networks) in CDSR, respectively, as shown in Figure 6. Moreover, from Figure 5, we can observe that the three basic technologies are not independent entities and can be applied interchangeably or in parallel based on the characteristics of different technologies. Notice that multi-layer perceptron (MLP) is used in almost all existing models for non-linear feature mapping or to simply aggregate cross-domain information Xu et al. (2023b); Hong and Jung (2023), so we do not describe it separately.
@@ 4.1.1 Recurrent Neural Networks in CDSR
Recurrent neural networks (RNNs) model sequences, by taking the current time step’s representation and the hidden features from the previous time step as input, and generating the output and hidden features for the current time step. Due to the issues of vanishing and exploding gradients, its variants, i.e., gated recurrent unit (GRU) and long short-term memory (LSTM) Yang et al. (2020), are more widely adopted. In CDSR, some researchers Li et al. (2021); Chen et al. (2021); Zang et al. (2023) directly employ GRUs as encoders of sequences to get sequential representations. While some researchers incorporate a shared unit into each step of RNNs to transfer cross-domain information. The shared unit could be a shared-account filter unit Ma et al. (2019); Sun et al. (2023) or the common representations of the overlapping users Wang et al. (2020). While RNNs are easy to implement and are capable of modeling the temporal relationships in sequences, they suffer from issues such as vanishing and exploding gradients during training.
@@ 4.2 Auxiliary Learning
In addition to the aforementioned key technologies, the utilization of auxiliary learning technologies to facilitate the integration of cross-domain information garner significant attention among researchers.
@@ 4.2.1 Transfer Learning
Transfer learning is primarily employed to transfer knowledge learned from one task to another task. For instance, CDNST Zhuang et al. (2020) transfers the novelty-seeking trait learned from a source domain to a target domain. SATLR Liu and Zhu (2021) considers transferring knowledge by multiplying the independently learned representations from one domain with an orthogonal mapping matrix. To improve the ability of knowledge transfer, some researchers propose dual transfer learning. DAT-MDI Chen et al. (2021) combines a dual transfer model with slot attention to self-adapte item embedding from different domains. DASL Li et al. (2021) applies a dual embedding component to unify the learning process of user representations and then proposes a dual attention component to incorporate user behaviors in multiple domains.
@@ 4.2.2 Contrastive Learning
Contrastive learning (CL) is also a widely applied technique that leverages the similarities and differences between samples to extract useful information. In the case of Tri-CDR Ma et al. (2023), it designs two contrastive learning tasks, i.e., the coarse-grained similarity modeling and the fine-grained distinction modeling. The coarse-grained similarity modeling closes three domains’ sequence representations of the same user, and the fine-grained distinction modeling assumes the distance between domain A and domain B should be larger than the distance between domain B and the hybrid domain. MGCL Xu et al. (2023c) views the local and global item representations of a user as the positive samples and the representations from different users as the negative samples. Additionally, some methods aggregate the sequences and then combine those processed sequences with CL. DREAM Ye et al. (2023) proposes supervised contrastive learning to minimize the relevance among inter-sequences with different preferences.
@@ 4.2.4 Discussion
Incorporating auxiliary learning technologies aims to enhance the model’s ability to capture cross-domain information. Contrastive learning can make learned representations more discriminative and robust, but it is sensitive to the designed contrastive strategy. Transfer learning enables the sharing of information across domains, but the effectiveness of knowledge transfer is affected by the correlation between domains. There is also adversarial learning that can unify representations from different domains but the training process is more complex.
@@ 6 Future Directions
Although certain progress has been made in addressing the CDSR problem, there is still ample room for improvement. In this section, we summarize several promising directions for potential developments from different perspectives.
@@ 7 Conlusions
Cross-domain sequential recommendation (CDSR) is an emerging problem that extends traditional recommender systems by incorporating both sequential and cross-domain granularities, offering the potential to effectively address data sparsity. In this paper, we approach CDSR from the perspective of a four-dimensional tensor and then provide a comprehensive overview of CDSR from both macro and micro views. From a macro view, we summarize the existing models by abstracting multi-level fusion structures. From a micro view, we focus on analyzing and summarizing the basic technologies and auxiliary learning technologies employed by the models. Finally, we also include some public datasets and experimental results for CDSR, and provide some promising future research directions.

### 19 Jan 2024 v2

@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
Below, we take a more tangible perspective to summarize the technologies adopted by existing models in addressing the challenges. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively.
@@ 4.1 Basic Technologies
According to Section 2.3, the cross-domain information is always considered as a pure sequential representation or a graph-encoded sequential representation. So we illustrate the utilization of sequence modeling technologies (i.e., recurrent neural networks and attention mechanisms) and graph structure modeling technologies (i.e., graph neural networks) in CDSR, respectively, as shown in Figure 6. Moreover, from Figure 5, we can observe that the three basic technologies are not independent entities and can be applied interchangeably or in parallel based on the characteristics of different technologies. Notice that multi-layer perceptron (MLP) is used in almost all existing models for non-linear feature mapping or to simply aggregate cross-domain information Xu et al. (2023b); Hong and Jung (2023), so we do not describe it separately.
@@ 4.1.1 Recurrent Neural Networks in CDSR
Recurrent neural networks (RNNs) model sequences, by taking the current time step’s representation and the hidden features from the previous time step as input, and generating the output and hidden features for the current time step. Due to the issues of vanishing and exploding gradients, its variants, i.e., gated recurrent unit (GRU) and long short-term memory (LSTM) Yang et al. (2020), are more widely adopted. In CDSR, some works Li et al. (2021); Chen et al. (2021); Zang et al. (2023) directly employ GRUs as encoders of sequences to get sequential representations. While some works incorporate a shared unit into each step of RNNs to transfer cross-domain information. The shared unit could be a shared-account filter unit Ma et al. (2019); Sun et al. (2023) or the common representations of the overlapping users Wang et al. (2020). While RNNs are easy to implement and are capable of modeling the temporal relationships in sequences, they still suffer from issues such as vanishing and exploding gradients during training.
@@ 4.2 Auxiliary Learning
In addition to the aforementioned key technologies, the utilization of auxiliary learning technologies to facilitate the integration of cross-domain information garners significant attention.
@@ 4.2.1 Transfer Learning
Transfer learning is primarily employed to transfer knowledge learned from one task to another task. For instance, CDNST Zhuang et al. (2020) transfers the novelty-seeking trait learned from a source domain to a target domain. SATLR Liu and Zhu (2021) considers transferring knowledge by multiplying the independently learned representations from one domain with an orthogonal mapping matrix. To improve the ability of knowledge transfer, some works adopt dual transfer learning. DAT-MDI Chen et al. (2021) combines a dual transfer model with slot attention to self-adapte item embedding from different domains. DASL Li et al. (2021) applies a dual embedding component to unify the learning process of user representations and then proposes a dual attention component to incorporate user behaviors in multiple domains.
@@ 4.2.2 Contrastive Learning
Contrastive learning (CL) is also a widely applied technique that leverages the similarities and differences between samples to extract useful information. In the case of Tri-CDR Ma et al. (2023), it designs two contrastive learning tasks, i.e., the coarse-grained similarity modeling and the fine-grained distinction modeling. The coarse-grained similarity modeling closes three domains’ sequence representations of the same user, and the fine-grained distinction modeling assumes the distance between domain A and domain B should be larger than the distance between domain B and the hybrid domain. MGCL Xu et al. (2023c) views the local and global item representations of a user as the positive samples and the representations from different users as the negative samples. Additionally, some works aggregate the sequences and then combine those processed sequences with CL. DREAM Ye et al. (2023) proposes supervised contrastive learning to minimize the relevance among inter-sequences with different preferences.
@@ 4.2.4 Discussion
Incorporating auxiliary learning technologies aims to enhance the model’s ability to capture cross-domain information. Contrastive learning can make learned representations more discriminative and robust, but it is sensitive to the designed contrastive strategy. Transfer learning enables the sharing of information across domains, but the effectiveness of knowledge transfer is affected by the correlation between domains, i.e., there is vulnerability to negative transfer. There is also adversarial learning that can unify representations from different domains but the training process is more complex.
@@ 6 Future Directions
Although certain progress has been made in addressing the CDSR problem, there is still ample room for improvement. In this section, we provide several promising directions in CDSR for potential developments from different aspects.
@@ 7 Conlusions
Cross-domain sequential recommendation (CDSR) extends traditional recommender systems by incorporating sequential and cross-domain information, thus addressing data sparsity. In this survey, we approach CDSR with a four-dimensional tensor and offer a comprehensive overview from macro and micro views. From a macro view, we summarize the existing models by abstracting multi-level fusion structures. From a micro view, we focus on analyzing and summarizing the basic technologies and auxiliary learning technologies employed by the models. Finally, we also include some public datasets and experimental results for CDSR, and provide some promising future research directions.